Gregory wants to maximise his well-being and has a marginal benefit curve (MB) of donuts represented by MB = 5 – 2q. If the market price of donuts is $6, how many donuts should Gregory buy?

a) 0

b) 15

c) 40

d) 10

e) 20
Answer #1

a) 0

(Well being is maximized at the point where P = MB
So, 6 = 5 - 2q
So, 2q = 5 - 6 = -1
So, q = -1/2
Thus, Gregory shouldn't buy any donuts because q is negative at the given price.)